# Krylov exponents and power spectra for maximal quantum chaos: an EFT approach

### v2

@@ Appendix A Recursion method for calculating Lanczos coefficients
A method to calculate Lanczos coefficients through a simple recursion relation is given in Dymarsky:2019elm . Given a thermal correlator $C(\tau)$, with $\tau$ Euclidean time, and regulated ($C(\tau):=\operatorname{Tr}{(\rho_{\beta}^{1/2}e^{-H\tau}V(0)e^{H\tau}\rho_{\beta}^{1/2}V(0)}$), one can find the Lanczos coefficients by solving the recursion relation
@@ Appendix B Physical properties of thermal autocorrelators
Here we summarise key physical properties of thermal autocorrelators. For more on these properties, as well as general background on thermal field theory, see the textbooks Bellac:2011kqa ; Das:1997gg ; Laine:2016hma ; Kapusta:2006pm .
@@ C.1 Defining the Meijer G-function
The Meijer G-function (see chapter 9.3 of gradshteyn2014table , sec. 5.3 in volume I of erdelyi1953htf1 or chapter 16 of NIST:DLMF ) can be defined through its Mellin-Barnes representation
@@
where $D=x\frac{\mathrm{d}}{\mathrm{d}x}$. The Meijer function is related to hypergeometric functions through Slater’s theorem; eq. (16.17.2) in NIST:DLMF ;
@@
see gradshteyn2014table Table 17.23, instance $20$, and
@@
for $x,y\in\mathbb{R}$, see eq. (8.332.4) in gradshteyn2014table .

### v3

@@ Appendix A Recursion method for calculating Lanczos coefficients
A method to calculate Lanczos coefficients through a simple recursion relation is given in Dymarsky and Gorsky (2020). Given a thermal correlator $C(\tau)$, with $\tau$ Euclidean time, and regulated ($C(\tau):=\operatorname{Tr}{(\rho_{\beta}^{1/2}e^{-H\tau}V(0)e^{H\tau}\rho_{\beta}^{1/2}V(0)}$), one can find the Lanczos coefficients by solving the recursion relation
@@ Appendix B Physical properties of thermal autocorrelators
Here we summarise key physical properties of thermal autocorrelators. For more on these properties, as well as general background on thermal field theory, see the textbooks Bellac (2011); Das (1997); Laine and Vuorinen (2016); Kapusta and Gale (2011).
@@ C.1 Defining the Meijer G-function
The Meijer G-function (see chapter 9.3 of Gradshteyn and Ryzhik (2014), sec. 5.3 in volume I of Erdélyi et al. (1981) or chapter 16 of Olver et al. (2010)) can be defined through its Mellin-Barnes representation
@@
where $D=x\frac{\mathrm{d}}{\mathrm{d}x}$. The Meijer function is related to hypergeometric functions through Slater’s theorem; eq. (16.17.2) in Olver et al. (2010);
@@
see Gradshteyn and Ryzhik (2014) Table 17.23, instance $20$, and
@@
for $x,y\in\mathbb{R}$, see eq. (8.332.4) in Gradshteyn and Ryzhik (2014).